# KunServe: Efficient Parameter-centric Memory Management for LLM Serving

### v4

@@ Introduction
Figure 14: An ablation study of running KunServe on Qwen-2.5-14B on LongBench dataset.
To study the effectiveness of each techniques proposed in §4, we conducted an ablation study on the system performance with different techniques incrementally enabled. Figure 14 shows the detailed study results on the LongBench dataset with Qwen-2.5-14B model. We omit other workloads and models due to space limitation since they have similar results. We report the end-to-end request latencies during the burst period in Figure 12.
Effectiveness of dyanmic parameter drop. First, we can see that parameter drop contributes (+Dynamic drop) to the most tail latency reductions. On the LongBench workload, the P90, P99 and P999 TTFT of KunServe are reduced by 8.8 $\times$, 11.7 $\times$ and 10.3 $\times$ compared to vLLM (DP). The key reason is that it completely eliminates queuing delays. Specifically, under bursts, there are 87 queued requests (whose TTFT > SLO(5 $\times$)) in this evaluation, KunServe executes them with enlarged GPU memory freed by dropping parameters. Though a larger batch size and pipeline bubbles lead to a TPOT increase in request processing (21–31.9% increase compared to the origin DP scheduling), it is still orders of magnitude smaller than queuing introduced by insufficient memory of vanilla vLLM.
@@
Effectiveness of lookahead batch formulation. Finally, with lookahead batch formulation (+ Lookahead), we further reduce the P90, P99 and P999 TPOT by 4.5%, 10.6% and 9.7%, respectively. The reduction in latencies directly comes from the more efficient pipeline execution. For example, without lookahead batch formulation, KunServe suffers 21.9% bubble time (defined as the ratio of idle GPU cycles) on average during the pipelined execution, while lookahead batch formulation effectively reduces the bubbles to 8.3%. This further results in an 20% throughput increase during the execution.
Figure 15: The performance of KunServe and its baselines in a long run (640s) of BurstGPT.
@@
While most of the techniques are designed to mitigate the queuing problem under memory throttling, KunServe also needs to restore the dropped parameters to keep the system running with the lowest latency under normal periods. To show the effectiveness of dynamic parameter restoration, Figure 15 presents the serving performance over a long run of BurstGPT workload with multiple memory throttling periods. To help understand the behavior of KunServe, we mark the time periods with dropping as grey boxes, other periods are running without parameter drop for KunServe.
First, we can see that dynamic restoration parameters when the memory pressure is low reduces the P50 latencies of TTFT and TPOT by 28 % and 23 %, respectively. This comes from the fact that with restoration, KunServe executes with the communication of pipeline parallelism. Without it, KunServe has 1.7 $\times$ higher TTFT compared with vLLM for the same reason.
Second, restoration even improves the P99 TTFT and TPOT by 6.4 $\times$ and 1.2 $\times$, respectively, though both approaches can free sufficient memory to avoid queuing. This is because without restoration, KunServe falls back to vLLM (PP) so it has a lower throughput under normal periods. As a result, KunServe w/o restoration suffers from larger bursts, as illustrated at the beginning of the second wave in the third row of Figure 15 (time 440s).
Performance under unrealistic extreme bursts
Figure 16: An evaluation of KunServe running Qwen-2.5-72B under extreme bursts.
While KunServe drops parameters to mitigate queuing, the memory that can be freed is bounded by the model size (see Table 1), so we have a limit in handling throttling caused by bursts. Nevertheless, KunServe can handle bursts much longer than existing systems, i.e., longer than any burst we have seen in the BurstGPT trace.
@@
Handling memory throttling with lossy methods. One possible way to handle memory throttling is to reduce the memory footprint of the serving, e.g., by compressing the activations [17, 28]. For example, FP8 quantization [44] reduces the token memory usage by 2 $\times$. And further, methods like SparseGPT [19] prune parameters to 50% sparsity. Unfortunately, such methods are lossy and can lead to model accuracy degradation or compromised user experience [30]. KunServe copes with the performance degradation caused by memory throttling without sacrificing the model accuracy.
Handling memory throttling with lossless methods. KunServe continues the line of work on handling memory throttling during LLM serving without modifying the model [27, 46, 37, 41, 50, 27, 5]. These works focus on allowing queued requests to execute by reorganizing GPU memory either with swap or migration-based methods, which does not create more space for execution so they either sacrifice ongoing requests or queued requests, as we have extensively analyzed in §2.3.
LLM serving optimizations. Considerable research has focused on improving the efficiency of LLM serving under abundant memory [16, 15, 7, 24, 27, 36, 50, 35]. KunServe builds on these works and seamlessly integrates with them. A recent work—POD-ATTENTION [26]—proposes a better chunked prefill implementations. It is orthogonal to our work and KunServe can benefit from its high performance kernel to get better performance in all states. NanoFlow [51] provides us with a more efficient microbatch scheduling, which is of help to KunServe after parameter dropping.
OS techniques for handling memory throttling. Handling memory throttling has been studied in operating systems for decades: e.g., Linux adopted a swap-based mechanism to handle memory pressure [1]. KunServe leverages the domain-specific knowledge of LLM serving to expose more memory to serving requests beyond the limit of a general-purpose swap-based method.
@@ Conclusion
In this paper, we are the first to demonstrate that parameter-centric memory management can effectively address the latency spikes caused by memory throttling in LLM serving. We built KunServe, an LLM serving system that cooperatively drops parameters to free up memory to eliminate all queuing under throttling. We also proposed a set of techniques to ensure all requests execute efficiently after parameter dropping, including drop plan generation with local unified memory management, coordinated KVCache exchange and lookahead batch formulation. KunServe reduces tail TTFT by up to 72.2 $\times$ compared to state-of-the-art systems like Llumnix, vLLM and InferCept.
@@ Acknowledgement
We sincerely thank Hanze Zhang, Tianle Sun, Wenxin Zheng, Jinrong Yang, Zhe Jiang, Zihang Zhong, Mingcong Han, Xian Xu, Yu Xia, Yingyi Hao, Hongrui Xie from IPADS for their valuable advice on this paper. We thank Mingyu Li from CAS for his advice. We also thank Wencong Xiao for his suggestions on paper writing. We thank Bytedance Seed Foundation team for their platform support.
Appendix A Appendix
Minimizing execution bubbles under bursts
Figure 17: (a) An illustration of the imbalanced batch execution time of existing chunking method. (b) How lookahead batch formulation balance microbatches via better token chunking. Inst. is abbreviation for instance.
We formally modeled the lookahead batch formulation problem to demonstrate its NP-hard nature. Most symbol definitions related to the modeling can be found in §4.3. We introduce the following symbols for clarity:
  • $\text{tokens}_{\_}b$: the token number of (micro) batch $b$.
  • $\delta_{\_}k$: the upperbound of the execution time of microbatch $b_{\_}{k+1}$. Typically, it is equivalent to $\max\left(\text{cost}_{\_}{b_{\_}k},\text{cost}_{\_}{b_{\_}{k+1}}\right)$ in minimizing process.
  • $\text{cost}_{\_}{upper}$: the upperbound of each microbatch in batch formulation.
Figure 17 illustrates how token-count-based chunking casuing bubbles and lengthens pipeline exectuion time. As shown in figure, for a pipeline formed by $m$ microbatches, the pipeline exectuion time (cost) can be calculated by:
$\text{cost}_{\_}{\text{pp}\left(\mathcal{B}\right)}=\text{cost}_{\_}{b_{\_}1}+\sum_{\_}{k=1}^{m-1}\max\left(\text{cost}_{\_}{b_{\_}k},\text{cost}_{\_}{b_{\_}{k+1}}\right)+\text{cost}_{\_}{b_{\_}m}$ (4)
By providing different minimal token constraint (i.e., $MIN$ in Figure 11), user can calculated $m$ as $\lceil\frac{\text{tokens}_{\_}\mathcal{B}}{MIN}\rceil$. For any given $m$, to distribute load evenly, we set the $\text{cost}_{\_}{upper}$ as $\lceil\frac{\text{cost}_{\_}{\mathcal{B}}}{m}\rceil$. The target of optimization is minimizing execution cost (see Figure 10) of $\mathcal{B}$. We have:
min $\text{cost}_{\_}{b_{\_}1}+\sum_{\_}{k=1}^{m-1}\delta_{\_}k+\text{cost}_{\_}{b_{\_}m}$
s.t. $\delta_{\_}k\geq\text{cost}_{\_}{b_{\_}{k+1}},\quad\forall k=1,\dots,m-1$
$\delta_{\_}k\geq\text{cost}_{\_}{b_{\_}k},\quad\forall k=1,\dots,m-1$
$\text{cost}_{\_}{b_{\_}k}\leq\text{cost}_{\_}{upper},\quad\forall k=1,\dots,m$
variables ($\text{variables }\left(\delta_{\_}k,c_{\_}{ij},x_{\_}{ij}\right)\text{ are integer}.$ (5)
Due to quadratic nature of cost function, the above problem is a polynomial programming. While a brute-force search is feasible when the request set $\mathcal{R}$ is small, the NP-hard complexity makes this approach impractical for online scheduling as the number of requests increases. In such case, our divide-and-conquer algorithm (Figure 10) plots a nearly-optimal execution plan quickly.

### v5

@@ Introduction
Figure 14. An ablation study of running KunServe on Qwen-2.5-14B on LongBench dataset. A smaller bubble time directly implies a better GPU utilization.
To study the effectiveness of each technique proposed in §4, we conducted an ablation study on the system performance with different techniques incrementally enabled. Figure 14 shows the detailed study results on the LongBench dataset with Qwen-2.5-14B model. We omit other workloads and models due to space limitation since they have similar results. We report the end-to-end request latencies during the burst period in Figure 12.
Effectiveness of dynamic parameter drop.    First, we can see that parameter drop contributes (+Dynamic drop) to the most tail latency reductions. On the LongBench workload, the P90, P99 and P999 TTFT of KunServe are reduced by 8.8 $\times$, 11.7 $\times$ and 10.3 $\times$ compared to vLLM (DP). The key reason is that it completely eliminates queuing delays. Specifically, under bursts, there are 87 queued requests (whose TTFT > SLO(5 $\times$)) in this evaluation, KunServe executes them with enlarged GPU memory freed by dropping parameters. Though a larger batch size and pipeline bubbles lead to a TPOT increase in request processing (21–31.9% increase compared to the original DP scheduling), it is still orders of magnitude smaller than queuing introduced by insufficient memory of vanilla vLLM.
@@
Effectiveness of lookahead batch formulation.    With lookahead batch formulation (+ Lookahead), we further reduce the P90, P99, and P999 TPOT by 4.5%, 10.6%, and 9.7%, respectively. The reduction in latency directly comes from the more efficient pipeline execution: without lookahead batch formulation, KunServe suffers 21.9% bubble time (the ratio of idle GPU cycles) on average during pipelined execution, while with it, the bubble time is only 8.3%. The reduced bubble time further improves throughput by 20%.
Figure 15. A comparison of execution latency estimated with our cost model and the real execution time of a Qwen-2.5-14B model in A800 GPUs. Left: the execution without prefix attention while right: the execution with prefix attention.
Accuracy of the batch formulation cost model
To evaluate the accuracy of KunServe’s cost model described in §4.3, we compare it with a baseline cost model neglecting attention computation cost found in existing work (nanoflow, ) and the ground truth. To demonstrate the generality of our model in both prefill and chunked prefill, we evaluate both requests without attention chunk ($R4^{0}$ in Figure 9 (c)) and with it ($R4^{1}$). As shown in Figure 15, for both cases, our cost model shows less than 5% deviation while the current formulation without considering attention has up to 48% and 74% deviations for requests without and with prefix attention, respectively. This confirms the importance of considering the attention computation cost in the cost model.
@@
To show the effectiveness of dynamic parameter restoration, Figure 16 presents the serving performance over a long run of BurstGPT workload with multiple overloading periods. To help understand the behavior of KunServe, we mark the time periods with dropping as grey boxes, other periods are running without parameter drop.
First, we observe that dynamic parameter restoration reduces the P50 latencies of TTFT and TPOT by 28 % and 23 %, respectively, due to the reduction of unnecessary pipeline execution. Second, restoration improves the P99 TTFT and TPOT by 6.4 $\times$ and 1.2 $\times$, respectively. Without restoration, KunServe falls back to vLLM (PP), resulting in lower throughput during normal periods, and consequently suffers from larger bursts with insufficient memory even with the dropped parameters, as illustrated at the beginning of the second wave in the third row of Figure 16 (time 440s).
Figure 16. The performance of KunServe and its baselines in a long run (640s) of BurstGPT.
Performance under extreme bursts
Figure 17. An evaluation of KunServe running Qwen-2.5-72B under extreme bursts.
While KunServe drops parameters to mitigate queuing, the memory that can be freed is bounded by the model size (see Table 1), so we have a limit in handling overloading caused by bursts. Nevertheless, KunServe can handle bursts much longer than existing systems, i.e., longer than any burst we have seen in the BurstGPT trace.
@@
Discussion
Supporting MoE models.    Our current implementation focuses on dense models, while Mixture of Experts (MoE) models are becoming increasingly popular recently: A key feature is that the inference of a request only activates a small subset of model parameters. For common MoE serving configurations like expert parallelism (EP) (DBLP:journals/corr/abs-2412-19437, ), KunServe seamlessly supports them because EP only changes the memory layout within an instance, while KunServe focuses on managing the GPU memory across serving instances. More importantly, KunServe is still effective for MoE models because KunServe only relies on the assumption that the model weights occupy a large portion of the GPU memory on an instance, which holds even with a sparse activation of experts (see Table 1). The assumption holds because though a request only requires a small portion of the model parameters, an instance still needs to load all the (large) model parameters to handle batches of requests that may activate all the experts.
Compatibility with different parallelism.  KunServe is compatible with different parallelism in LLM serving. KunServe only changes the parameter layout across instances in layer granularity, which is orthogonal to both the intra-layer layout change (e.g., EP and TP) within one instance and instance cooperation in SP. Thanks to the LLM’s modular structure, the intra-instance and inter-instance parallelism techniques can be applied together (DBLP:journals/corr/abs-1909-08053, ; loongserve, ).
Comparison with autoscaling.    Autoscaling—adding more instances to handle overloads—is also a common approach to handle memory overloading (DBLP:conf/osdi/ZhangWLWS0025, ). A key difference is that KunServe does not have the cold start time—the time to make an instance capable of serving. Thus, KunServe is better than autoscaling in cases where dropping alone is able to handle the overloading, as the cold start time is typically non-trivial for LLM providers (DBLP:journals/corr/abs-2407-00079, ). Nevertheless, for long bursts (see §5.6), KunServe still incorporates autoscaling since the memory that can be freed by dropping is limited: the continuously coming requests from the burst will exhaust all the free memory freed by dropping.
@@ Discussion
Handling memory overloading with lossy methods.   One possible way to handle memory overloading is to reduce the memory footprint of the serving, e.g., by compressing the activations (llm-quant, ; quant-eval, ). For example, FP8 quantization (llm-compressor, ) reduces the token memory usage by 2 $\times$, and methods like SparseGPT (sparsegpt, ) prune parameters to 50% sparsity. Unfortunately, such methods are lossy and can lead to model accuracy degradation or compromised user experience (quant-damage, ). KunServe copes with the performance degradation caused by memory overloading without sacrificing the model accuracy.
Handling memory overloading with lossless methods.KunServe continues the line of work on handling memory overloading during LLM serving without modifying the model inference (vllm, ; DBLP:journals/corr/abs-2404-09526, ; DBLP:journals/corr/abs-2407-00079, ; llumnix, ; distserve, ; vllm, ; infercept, ). These works focus on allowing queued requests to execute by reorganizing GPU memory either with swap or migration-based methods, which do not create more space for execution so they either sacrifice ongoing requests or queued requests, as analyzed in §2.3. In contrast, KunServe frees more memory for execution with a new parameter-centric memory management method.
LLM serving optimizations.  Considerable research has focused on improving the efficiency of LLM serving under abundant memory (flashattention, ; flashattention2, ; sarathi-v1, ; fastgen, ; vllm, ; DBLP:journals/corr/abs-2405-04437, ; distserve, ; splitwise, ). KunServe builds on these works and seamlessly integrates with them. A recent work—POD-ATTENTION (podattn, )—proposes a better chunked prefill implementation. It is orthogonal to our work and KunServe can benefit from its high-performance kernel to get better performance in all states. NanoFlow (nanoflow, ) provides us with a more efficient microbatch scheduling, which is of help to KunServe after parameter dropping.
OS techniques for handling memory overloading.  Handling memory overloading has been studied in operating systems for decades: e.g., Linux adopted a swap-based mechanism to handle memory pressure (mglru, ). KunServe leverages the domain-specific knowledge of LLM serving to expose more memory to serving requests beyond the limit of a general-purpose swap-based method.
@@ Conclusion
In this paper, we are the first to demonstrate that parameter-centric memory management can effectively address the latency spikes caused by memory overloading in LLM serving. We built KunServe, an LLM serving system that cooperatively drops parameters to free up memory to eliminate queuing under overloading. We also proposed a set of techniques to ensure all requests execute efficiently after parameter dropping, including drop plan generation with local unified memory management, coordinated KVCache exchange and lookahead batch formulation. KunServe reduces tail TTFT by up to 72.2 $\times$ compared to state-of-the-art systems like Llumnix, vLLM and InferCept.
@@ Acknowledgement
We sincerely thank our shepherd Jayashree Mohan and the reviewers from OSDI’25 and EuroSys’26 for their insightful feedback. We are grateful to Wencong Xiao from ByteDance, Mingcong Han, Hanze Zhang, Xian Xu, Yu Xia, Yingyi Hao, and Hongrui Xie from IPADS for their valuable advice. We also thank the ByteDance seed-foundation team for their platform support. We thank Chao Fei from KAUST for his contributions to the codebase of KunServe. This work was supported in part by the National Natural Science Foundation of China (No. 62572302 and 62272291), and the Fundamental Research Funds for the Central Universities.